OpenStudy (anonymous):

At what y points on the curve xy = 25 is the tangent line parallel to 5x + y = 0?

OpenStudy (amistre64):

when its slope is -a/b given the form ax + by = c

OpenStudy (amistre64):

the derivative of xy=25 wrt x is: y + xy' = 0; y' = -y/x = -5/1

OpenStudy (amistre64):

need some x,y in xy that comes at to 25 and slopes to -5/1

OpenStudy (amistre64):

or something like that :)

OpenStudy (anonymous):

plus minus sqrt 5?

OpenStudy (amistre64):

we could look at this as y = 25/x to get y' = -25/x^2 = -5

OpenStudy (amistre64):

to wit estudier gives good advise

OpenStudy (anonymous):

-25/x^2 = -5